OpenStudy (anonymous):

ann,david, john are are at a coffee shop. ann orders four coffees, one juice and two donuts and pays$10.90. john orders one coffe and three donuts and pays $5.35.david orders three coffees one juice and four donuts and pays $11.80. find the price of coffee donuts and juice at this shop. A. set up a linear system for the prices of the coffee juice and donuts B. use cramers rule to solve the system remember to show all work

OpenStudy (anonymous):

Let coffee = x, juice = y, donuts= z Ann : 4x+y+2z=10.90 John : x+3z=5.35 David : 3x +y +4z = 11.80

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Then find D, Dx, Dy and Dz to solve for x,y,z

OpenStudy (anonymous):

i dont know how to solve for those

OpenStudy (anonymous):

\[\left(\begin{matrix}10.9 \\ 5.35\\ 11.8\end{matrix}\right)=\left(\begin{matrix}x \\ y\\z \end{matrix}\right)\left[\begin{matrix}4 & 1 &3 \\ 1 & 3 & 1\\ 2&0&4\end{matrix}\right]\]

OpenStudy (anonymous):

sorry it should be vertically 1st column 4,1,3 2nd column 1,31 and 3rd column 2,0,4

OpenStudy (anonymous):

http://www.purplemath.com/modules/cramers.htm